# A tweezer array with 6100 highly coherent atomic qubits

### v3

@@ Abstract
Finally, we demonstrate global coherent single-qubit operation of our atom array and probe single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023. To compensate for the inhomogeneous Rabi frequency across the array, which would result in pulse-length errors, we use the SCROFULOUS composite pulse cummins_tackling_2003. We apply gates sampled from the 24 unitaries composing the Clifford group , followed by an inverse operation, and measure the final population in (Fig. 4d). Fitting the decay as the number of gates increases yields an average Clifford gate fidelity , limited by phase noise in our system likely due to magnetic field noise (Methods). This could be readily addressed by upgrading the current sources driving the magnetic field coils, shielding the vacuum cell, or by operating at MHz-scale by driving optical Raman transitions levine_dispersive_2022.
@@
Towards quantum computing with a zone-based architecture — coherent transport
@@
We now focus more specifically on the practical implementation of a quantum computer, as it is a flagship application of our work and because it demands the most sophisticated toolbox of aforementioned use cases. Universal quantum computing requires single-qubit and two-qubit gates which can be implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
@@
Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds than grid-like connectivity bravyi_tradeoffs_2010, presenting a promising pathway to large-scale fault-tolerant quantum computation. A zone-based architecture also provides a path for mid-circuit readout bluvstein_logical_2024.
We depict a possible zone layout in Fig. 5a, which includes a storage zone large enough for more than 6,100 atoms. We do not foresee challenges in creating the zones themselves, e.g., Rydberg-based two-qubit gates should be feasible in a large interaction zone for more than 500 gates in parallel with state-of-the-art fidelities (SI §IV).
@@
However, coherence-preserving transport between storage and adjacent interaction or readout zones might require covering large distances of 500 m. While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is several times further than previously demonstrated single-atom transport bluvstein_quantum_2022, bluvstein_logical_2024. To address this challenge, we investigate the feasibility of coherent transport at this scale by loading atoms directly into tweezers made by AODs and performing the movement, given that pick-up and drop-off effects are not dependent on the distance of transport. We utilize the most wide-band commercially available AODs at near-infrared wavelengths, which for the optical parameters used here cover up to m along one axis (Methods).
In probing the survival of atoms as they are moved long-distance with differing speeds, we find that the speed of transport is strongly constrained by cylindrical lensing, an effect that occurs as the AOD frequency is rapidly swept dickson_optical_1972, and which is more deleterious as the AOD field of view is scaled up (SI). In Fig. 5b, we present data that illustrate a striking consequence of this effect: with the same trap depth and trajectory, atoms can be transported considerably faster in a diagonal direction, utilizing two crossed AODs, when compared with using one AOD only. Such diagonal moves effectively turn cylindrical lensing into spherical lensing, thereby suppressing the worst impact. In the SI, we describe how diagonal motion can be used to transport atoms between zones without any penalty compared with a usual configuration relying on straight moves. We also discuss other known optical strategies that could completely remove any lensing friedman_acousto-optic_2000.
With diagonal moves, we first demonstrate in Fig. 5c the coherence of atoms being held in static tweezers in comparison with the coherence of atoms being transported by 610 m in 1.6 ms. We suppress any dephasing with one XY4 dynamical decoupling sequence per move, and do not observe loss in coherence between the static and transported case.
@@
However, for any realistic application of coherent transport, atoms will have to be moved multiple times while maintaining coherence. Thus, we characterize the fidelity of the quantum channel defined by the coherent transport, through interleaved randomized benchmarking magesan_efficient_2012 (Fig. 5d). To our knowledge, this the first such rigorous transport channel fidelity characterization for neutral atoms. To maximize the dephasing cancellation, we apply dynamical decoupling in a transformed Clifford frame (Methods).
@@
We perform this benchmarking technique for an average move distance from the storage zone to the interaction zone of m (Fig. 5e,g), and also for a long move distance of m (Fig. 5f). For each set of parameters (Fig. 5e-g), we first characterize the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel). For selected durations, we then characterize the actual randomized benchmarking signal, the return probability to the initial quantum state (middle panel).
@@
Notably, the resulting return probability data do not follow an exponential trend, because at large numbers of moves they become dominated by losses from the trap; as such, the fidelity for the transport channel depends on the number of previously executed moves. We thus define an instantaneous fidelity, the fidelity of the transport channel after a certain number of previous moves, shown in the bottom panel (Methods).
The instantaneous fidelity approaches a constant value for small number of moves (), where losses are the sub-dominant error. This regime is most relevant for QEC, since one can in principle swap data qubits and ancilla qubits every few layers of gates chow_circuit-based_2024. We find that the fidelities in this regime approach 99.95% for both transport distances of 270 m (Fig. 5e) and 610 m (Fig. 5g) in ms and in ms respectively using mK deep traps. We note that there is a small discrepancy between the early-time instantaneous fidelities obtained for 270 m and 610 m, and anticipate it could be remedied through more careful trap depth calibration along the trajectory.
For some applications, one might desire to optimize on the speed of movement and use a deeper trap to do so. In Fig. 5g, we show that atoms can be moved by 270 m in 400 s with a trap depth of mK, at the cost of a reduced fidelity of 99.8%. Comparing Fig. 5e and Fig. 5g illustrates an important trade-off pertaining to coherent transport: while atoms can be moved faster by increasing the trap depth , the associated transport fidelity for small number of moves is also reduced. In the limit where noise is entirely induced by tweezer intensity fluctuations, this can be understood by noticing that the dephasing strength scales as when the required duration for long-distance transport merely scales as . We note that experimentally in static traps, we find an even stronger scaling of coherence time than linear in , likely due to other sources of noise (Ext. Data Fig. 7g).
@@
Finally, we envision utilizing multiple pairs of crossed AODs to cover the full extent of the array in Fig. 5a, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions. With the layout presented here, four such regions would be necessary. Alternatively, additional scanning techniques, e.g., fast-scanning mirrors, can be used to position the field of view of single pair of crossed AODs across the full array in an iterative fashion.
@@
As we lay out in detail in the SI, such techniques can be used also for initial rearrangement of atoms into the storage zone. For example, by implementing the recently developed “Tetris algorithm” wang_accelerating_2023 in four quadrants, with conservative estimates for relavant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in 137 ms or sequentially quadrant-by-quadrant in 522 ms. Further, this estimate is based on straight moves along a single AOD axis, which could be sped up by utilizing techniques for mitigating cylindrical lensing (SI §III.B).
@@
We have shown that while scaling qubit numbers in optical tweezers to more than 6,100, we can simultaneously achieve record-high imaging survival alongside competitive imaging fidelity. We measure record room-temperature vacuum-limited lifetime and coherence time in alkali-metal atom tweezer arrays, with a high global single-qubit gate fidelity that is limited by technical noise. We additionally characterize the fidelity of a long-distance quantum transport channel utilizing randomized benchmarking and providing clear evidence for the potential to use this large scale array with a zone-based architecture for quantum computing (SI).
@@
Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, as described in the previous section. Additionally, large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024 are made accessible through this work. For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2024 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here.
@@
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
@@
Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented by utilizing an objective with a housing material that retains less heat or with integrated cooling strategies. Such upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work (SI). Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023, norcia_iterative_2024_mod. Already in the near term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.
@@ Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Harry Levine, Zunqi Li, Ran Finkelstein, Pascal Scholl, Joonhee Choi, Dolev Bluvstein, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. H.J.M. acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. K.H.L. acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.Data availability

### v4

@@ Abstract
Finally, we determine single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023. To compensate for the inhomogeneous Rabi frequency across the array, we use the SCROFULOUS composite pulse cummins_tackling_2003. We apply gates sampled from the Clifford group , followed by an inverse operation, and measure the final population in (Fig. 4d). Fitting the decay as the number of gates increases yields an average Clifford gate fidelity , limited by phase noise in our system likely due to magnetic field noise (Methods). This could be readily addressed by upgrading the current sources driving the magnetic field coils or by operating at MHz-scale Rabi rates with optical Raman transitions levine_dispersive_2022.
@@
Towards zone-based quantum computing — coherent long-distance transport and atom transfer
@@
We now focus more specifically on the practical implementation of a quantum computer, as it is a flagship application of our work and because it demands the most sophisticated toolbox of aforementioned use cases.
Universal quantum computing requires single-qubit and two-qubit gates which have been implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
@@
Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds bravyi_tradeoffs_2010. This architecture also provides a path for mid-circuit readout bluvstein_logical_2024. We depict a possible zone layout in Ext. Data Fig. 10a and the Supplementary Information (SI), which includes a storage zone large enough for more than 6,100 atoms. We do not foresee challenges in creating the zones themselves, e.g., Rydberg-based two-qubit gates should be feasible in a large interaction zone for more than 500 gates in parallel with state-of-the-art fidelities (SI §IV).
@@
However, coherence-preserving transport between storage and adjacent interaction or readout zones might require covering large distances of 500 m. While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is significantly further than previously demonstrated distances for single-atom transport with tweezers bluvstein_quantum_2022, bluvstein_logical_2024. Furthermore, transferring atoms between dynamic (AOD-generated) and static (SLM-generated) tweezers requires precise relative alignment, conceivably challenging in our system due to the high laser power or potential for worsening aberrations over the large FOV.
Thus, we investigate the feasibility of coherence-preserving transport and SLM-AOD trap transfer over larger length scales. First, isolating challenges with the coherent transport operation, we load atoms directly into ten AOD-generated tweezers and characterize coherent moves up to 610 m (Fig. 5, top section of Ext. Data Fig. 10). Second, we assess the viability of large-scale parallel AOD-SLM trap transfers with 195 AOD-generated tweezers that span a square of 504 m 468 m (Fig. 6). As an outlook, we demonstrate a proof-of-principle combination of these techniques in a large-scale static array (although in a different trap layout) moving a 2D array of 47 atoms over 375 m, a distance comparable to predicted zone spacings in our system (Ext. Data Fig. 10e, f). For all operations, we utilize the most wide-band commercially available AODs at near-infrared wavelengths, which cover up to m along one axis for the optical parameters used here (Methods).
In probing atom survival as a function of long-distance movement speeds, we find that the speed of transport is strongly constrained by cylindrical lensing—an effect that occurs when the AOD frequency is rapidly swept dickson_optical_1972—which becomes increasingly deleterious as the AOD field of view is increased (SI III.A). Notably, using a pair of crossed AODs for diagonal transport converts cylindrical lensing into spherical lensing, enabling significantly faster movement (Fig. 5a). With diagonal moves, we first demonstrate in Fig. 5b negligible loss of coherence for atoms transported by 610 m in 1.6 ms. We suppress dephasing with one XY4 dynamical decoupling sequence per move.
@@
Realistic applications of coherent transport involve multiple consecutive moves. Therefore, we characterize the fidelity of the quantum channel defined by coherent transport through interleaved randomized benchmarking magesan_efficient_2012 (IRB, Fig. 5c, Methods). To the best of our knowledge, such a quantitative characterization of transport fidelity in neutral atoms has not been previously demonstrated. To maximize the dephasing cancellation, we apply dynamical decoupling in a transformed Clifford frame (Methods).
@@
We perform this benchmarking technique for a distance of m (Fig. 5d), with diagonal moves. We first measure the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel). For a 1.6-ms move using mK-deep traps, we then characterize the return probability to the initial quantum state after IRB as a function of the number of moves (middle panel). Other distances, trap depths and move times are shown in Ext. Data Fig. 10.
@@
The resulting IRB return probability data is non-exponential in the number of moves, because at large numbers of moves, trap losses become dominant and the fidelity for the transport channel depends on the number of previously executed moves. This motivates defining an instantaneous fidelity; i.e., the fidelity of the transport channel after a certain number of previous moves (Methods), shown in the bottom panel of Fig. 5d. The instantaneous fidelity approaches a constant value of for small numbers of one-way moves (), where losses are the sub-dominant error. This regime is most relevant for QEC, since data qubits and ancilla qubits can, in principle, be swapped every few layers of gates chow_circuit-based_2024_new.
We then move on to characterizing the atom transfer between static and dynamic tweezers. We demonstrate that these operations proceed without the emergence of unexpected technical challenges by performing high-fidelity parallel AOD-SLM transfer across the full field of view of the AOD (Fig. 6).
We use 195 AOD tweezers spread across 504 m 468 m (Fig. 6a) to perform and characterize the repeated transfer procedure, post-selected on initially filled SLM sites. As with coherent transport benchmarking, we evaluate the transfer fidelity as a function of the number of one-way transfers through IRB (Fig. 6d). To execute faster (or higher-fidelity transfers at a given duration), we propose and implement a trajectory where AOD ramp-and-move operations are simultaneously optimized with machine learning techniques to maximize survival (Fig. 6c, Methods). Compared with our manually optimized trajectory (Fig. 6b), this technique yields significantly higher atomic survival, and enables a one-way transfer fidelity of 99.81(3)% for transfers (compare blue and gray data in Fig. 6d, top panel).
In the future, such machine learning techniques could also be used to optimize combined pick-up and transport, where we find a fidelity of for the first operations at the chosen timescales with manually optimized methods (Methods, Ext. Data Fig. 10f).
@@
Finally, to cover the full extent of the array, we envision utilizing multiple pairs of crossed AODs, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions (SI Fig. 2). With the layout presented in Ext. Data Fig. 10a and the SI, four such regions would be necessary. Alternatively, additional scanning techniques (e.g., fast-scanning mirrors) can be used to position the field of view of a single pair of crossed AODs across the full array iteratively.
@@
Such techniques are also applicable to initial rearrangement of atoms in the storage zone. For example, by implementing a parallel assembly algorithmtian_parallel_2023, wang_accelerating_2023 in four quadrants (SI §II), with estimates for relevant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in 137 ms or sequentially quadrant-by-quadrant in 522 ms.
@@
We have shown scaling of neutral atom qubit numbers in optical tweezers to more than 6,100. We simultaneously achieve high imaging survival and fidelity as well as a long room-temperature vacuum-limited lifetime. We find record coherence times in alkali-metal atom tweezer arrays and a high global single-qubit gate fidelity, limited by technical noise. Further, we additionally characterize the fidelity of quantum transport channels for moves and trap transfer at relevant length scales, utilizing randomized benchmarking.
@@
Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, particularly relevant for QEC bravyi_high-threshold_2024, xu_constant-overhead_2024. Additionally, large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024 are made accessible through this work. For example, our platform—with the demonstrated qubit numbers—could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2025. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2024 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here.
@@
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (reducing the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
@@
Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented by utilizing an objective with a housing material that retains less heat or with integrated cooling strategies. Such upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work. Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023, norcia_iterative_2024_mod. Already in the near term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.
@@ Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Harry Levine, Zunqi Li, Ran Finkelstein, Pascal Scholl, Joonhee Choi, Dolev Bluvstein, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044), and the Heising-Simons Foundation (2024-4852). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. H.J.M. acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. K.H.L. acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.Data availability